Find ∫f(x)dx If, f[(3x-4)/3x+4)] = x+2

Find ∫f(x)dx


If,


f[(3x-4)/3x+4)] = x+2

Grade:11

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Shreyans

first calculate f(x)

put 3x-4/3x+4 = y

calculate x in terms of y

x=4(1+y)/3(1-y)

x+2=2(5-y)/3(1-y) = 2(y-5)/3(y-1)

f(y)=2(y-5)/3(y-1)

f(x)=2(x-5)/3(x-1)

to integrate this use partial fractions

f(x)=2/3 [(x-1-4)/(x-1)]= 2/3[1 - 4/x-1]

now integration is simple

∫2/3[1 - 4/x-1]dx

2/3[x-4log(x-1)] + C

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 


 

 

yash dalmia
13 years ago

put 3x-4/3x+4=t,

on solving by componendo and dividendo,

6x/-8=t+1/t-1

x=4(t+1)/3(1-t)

put this in the given condition, get f(x), then integrate it

Kushagra Madhukar
3 years ago
Dear student,
Please fidn the solution to your problem.
 
first calculate f(x)
put 3x-4/3x+4 = y
calculate x in terms of y
x=4(1+y)/3(1-y)
x+2=2(5-y)/3(1-y) = 2(y-5)/3(y-1)
f(y)=2(y-5)/3(y-1)
f(x)=2(x-5)/3(x-1)
to integrate this use partial fractions
f(x)=2/3 [(x-1-4)/(x-1)]= 2/3[1 - 4/x-1]
now integration is simple
∫2/3[1 - 4/x-1]dx
= 2/3[x-4log(x-1)] + C
 
Thanks and regards,
Kushagra